Prev Next

SW&IT Meeting Minutes 2020-02-10

Attendees: Carl, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200127
LR20200127 - Bec, Matthieu 80.0
LR20200127 - Holmberg, Carl 80.0
LR20200127 - Edmeades, David 72.0
LR20200127 - Sola, Igor 40.0
LR20200127 - Kubanek, Petr 80.0
LR20200127 - Hooper, Stephen 80.0
LR20200127 - Zhang, Yang 24.0
LR20200127 - Zhang, Xianyu 40.0

News from management

  • AO lead interviews
  • S/W posts moving into the new system

Previous Action Items

Owner Action Due DateSorted ascending Completion Date
DE Summit Veam Backup (soul-telemetry, etc) + bring back VM2 until PEPSI/VM arrives 2020-01-20  
IS Luci passwords in google doc 2020-02-10  

Open topics

  • IRTC server on obs1 (rpm)
  • aluminizing PC check
  • NAS status
  • IssueTrak, rm507-1/x e.g., handling IT8050
  • umac s/w contact for M2 hexapod

Team reports

MB

  • Last week
    • OC, WPC, Management
    • AHS on hydrogen/ioc.mountain
    • TMS kickoff
    • sciarc zoominar
  • This week
    • RemoteOps kickoff
    • AR TMS user stories kickoff
    • Argos cal source and alarms follow-up
    • masking alarms in ALH e.g. DX adsec, but others
    • expresso machine sticker
    • safety courses
    • TCS ITs review with PK and YZ
    • DMS developments

PK

  • 2020B patch
  • PCS patch for TTL constructors
  • SX? M3 support
  • DX M2 hexapod support
  • almost finished Malaga paper

DE

  • MPB refreshed and ready for assignment
    • UA hazmat team will collect swollen LiIon batteries
  • Summit emergency laptop arrived, configuring routes
  • Computer for Alexander Becker ordered; long lead time from Lenovo
  • Tucson Synology activities
    • 2 drives showing remapped sectors; swapping one at a time
    • Unresponsive Thurs evening, logs not illuminating
    • Reboot caused primary unit swap, rebuilding continued in background
    • Disk kicked from array on RS1 (now passive unit), unknown cause but it appears to have been transient. Rebuilding on original disk. Positive progress as of Fri afternoon.
    • Ordered spare 1TB SSD, holding off on swap until above rebuild is done to minimize load
    • Ordered spare 8TB drive; Seagate charges nonrefundable fee for advance replacement program; I'd rather spend money on a thing we get to keep.
  • Veeam licensing options investigate
    • I don't see a reason to upgrade at this time. Some nice features, but our needs are not complex and saving ~$3k seems like the logical choice
    • Adding a node license verified by Veeam rep to work with our existing install
  • OVIRT VM work stays on the list

XZ

  • Last week
    • Power Backplane meeting with Roberto
    • Send PBP meeting notes to Guido
    • Supoort SOUL
    • DX AdSec issue analysis and tests
    • DX AdSec reseating and swap with rigid AdSec
  • This week
    • Remote support DX AdSec test in the lab
    • Power Backplane?

YZ

  • Follow up with Olga on MODS file checksum script (cron job setup, may need to create new alert email alias).
  • Investigate LBC daily file moving script (script appeared to be in github, will speak with Igor and then plan to update it).
  • Some TCS and PCS test preparation.
  • TMS kick off meeting.
  • Investigate some TCS ECS GUI issues (instrument cooling alarm color issue).

SH

  • upgraded two rm507 machines
  • LBC KVM and LBC with Leroy
  • removed cron ssh/web rsync dependency from telemetry transportation
  • moved obs1 to local "/usr/local" rather than NFS shared "/usr/local".
  • Tucson Synology lock up hard reboot.
  • IRS exploration (brief)

next:
  • upgrade NUCs
  • re-kick hydrogen, name change
  • rm507 computer work
  • remote observing meeting?

CH

Last week

  • AGW: Write IT to fix AGW firmware address conflict.
  • AHS: HBS alert TCS changes deployed.
  • ARGOS: IT6224, initial discussion with Gustavo Rahmer re AOS data requirements.
  • Safety training.
  • Support scripts: updated RB_Science based on Stack Overflow feedback.
  • TCS: Shadowed PK for TCS 2020B code changes.

This week

  • AHS: follow up with LBTI re alert data collection.
  • ARGOS: IT6224, follow up discussion with Gustavo Rahmer, review code data entry points.
  • ARGOS: Review code for data entry points.
  • Support scripts: test updated RB_Science.

IS

Last week:
  • OT
    • OT bug and enhancement work. Two test releases out, looking good so far probably ready to build final release this week.
This week:
  • OT
    • Final release
  • PIT
    • Start working on PIT enhancements

  • Time Accounting Tool
    • Export to excel
    • Use django admin login capabilities for login and queue run program editing.
Topic revision: r16 - 10 Feb 2020, CarlHolmberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback